# Attosecond physics hidden in Cherenkov radiation

### v2

@@ Introduction
and it vanishes for classical emission with no recoil, $\omega/\varepsilon\to 0$. For materials like Al, Si, Be, and Ti, Cherenkov radiation can be observed in the soft X-ray range at the frequencies [39, 40, 38] $\omega\approx 72.5$, $100$, $110$, $453.8$ eV, respectively, which for $\varepsilon\sim(5-20)m$ yields $\Delta\theta_{\infty}<1-2\,\text{deg}$ (cf. Fig.3 c.), d.)).
Between the above points, the time $t_{d}$ becomes negative – see Fig.3 b.) – and the Gouy phase $\arctan t^{\prime}/t_{d}$ changes it sign, as if the electron packet shrinks during the emission. In the classical regime with no recoil, both the points merge and so $t_{d}$ and $L_{f}$ turn to infinity at the Cherenkov angle. Indeed, in a vicinity of this angle the spreading time with the recoil kept is
$\displaystyle t_{d}\Big{|}_{\cos\theta=1/nu_{p}}=\frac{2\varepsilon}{\sigma^{2}}\frac{n^{2}}{1-n^{2}}<0,$ (12)
where $\varepsilon=\gamma m$. We compare this with the spreading time of an electron packet, which is at rest on average in vacuum [41] $t_{d}^{(\text{e, rest})}=m/\sigma^{2}$. In the laboratory frame this time is $\gamma$ times larger, which coincides with $|t_{d}|$ up to the factor $22n^{2}/(n^{2}-1)>$ 2. So, in a vicinity of the Cherenkov direction spreading of the photon seems to reverse back and it is intimately connected with spreading of the electron packet.
@@
Dependence of the Wigner function (5) on the detection time $t$ comes exclusively from the following envelope:
@@
Here natural duration of the Cherenkov flash is defined by $\sigma_{t}(t^{\prime})$ and a time instant at which the probability to catch the photon around the point ${\bf r}$ is maximized is $t_{0}={\bm{l}}_{0}\cdot\left({\bf r}+(\partial_{{\bf p}}+\partial_{{\bf k}})\zeta_{fi}\right),\,{\bm{l}}_{0}=\left[({\bm{u}}_{p}-{\bm{u}}_{k})\times\left[{\bm{u}}_{k}\times{\bm{u}}_{p}\right]\right]/\left[{\bm{u}}_{p}\times{\bm{u}}_{k}\right]^{2}$.
One can neglect the term with the phase $\zeta_{fi}$ in ${\bm{R}}$ (6) in the wave zone where $r\to\infty$, and then the Wigner function (5) and the emitted energy are concentrated in a vicinity of the electron classical trajectory [47], ${\bf r}\sim{\bm{u}}_{p}t$. The detector registers a photon in the far field emitted at $t=0,r_{0}=0$ by a classical point-like electron at the time instant
@@
which will be called the classical arrival time. Let us compare this prediction with the above $t_{0}$, derived quantum mechanically. Directing the $z$ axis along the electron momentum ${\bf p}$, we find ${\bf k}=n\omega{\bm{l}},\ {\bm{l}}=\{\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta\},\ r\equiv{\bf r}\cdot{\bm{l}},\ t_{\text{cl.}}^{(\text{far-f.})}={\bf r}\cdot{\bm{l}}\,n,$ and ${\bm{l}}_{0}=\sin^{-2}\theta\,\left(\left(u_{p}^{-1}-n\cos{\theta}\right){\bf p}/|{\bf p}|+\left(n-u_{p}^{-1}\cos\theta\right){\bm{l}}\right)$. In a vicinity of the Cherenkov angle, we have ${\bm{l}}_{0}\to{\bm{l}}\,n$ and so for $\zeta_{fi}=0$ we get $t_{0}\to t_{\text{cl.}}^{(\text{far-f.})}={\bf r}\cdot{\bm{l}}\,n$, in accord with Eq.(15).
@@
However the phase $\zeta_{fi}$ cannot be ignored in the formation zone and this affects the photon arrival time $t_{0}$. In this regime, we call $t_{\text{cl.}}={\bf r}\cdot{\bm{l}}_{0}$ for arbitrary emission angles and so the quantum shift is
$\Delta t=t_{0}-t_{\text{cl.}}={\bm{l}}_{0}\cdot(\partial_{{\bf p}}+\partial_{{\bf k}})\zeta_{fi}.$ (16)
@@
The physical origin of this delay is the electric dipole moment density $\propto e(\partial_{{\bf p}}+\partial_{{\bf k}})\zeta_{fi}$ induced in medium by the field of the electron. We deal with an analogue of the AC Stark effect [16] with the atoms being off-resonantly polarized by a broadband spectrum $\Delta\omega$ of pseudo-photons. Similarly to the observed time delays when a laser propagates in a medium [48, 49, 50], here we encounter delays induced by the virtual photons, reemitted as real ones. Classically, one can look at this as if the photon was emitted not from a point-like electron, but from a point shifted laterally to $\Delta\rho\sim\beta\gamma\lambda/2\pi$ from its trajectory, which is a mean free path of the virtual photon [29, 12]. Numerically $\Delta\rho/c\sim\beta\gamma\lambda/2\pi c=\beta\gamma/\omega\sim 1\,\text{fs}-100\,\text{fs}$ for photons from IR to UV ranges and $\gamma=\varepsilon/m\sim 10$.
@@
Figure 4: The quantum shift $\Delta t=t_{0}-t_{\text{cl.}}={\bm{l}}_{0}\cdot(\partial_{{\bf p}}+\partial_{{\bf k}})\zeta_{fi}$ from Eq.(16) of the photon arrival time compared to the classical value and the Cherenkov flash duration (the green lines with $\sigma=p_{\perp}$). The electron energy is typical for a transmission electron microscope: $\beta=0.7$, and $p^{\prime}_{\perp}=0.99\times p_{\perp},p^{\prime}_{z}=0.9\times\beta,n=1.5$. The left panel (a.): $p_{\perp}=10^{-5}\,m,\,1/p_{\perp}\gtrsim 10\,\text{nm}$, the right panel (b.): $p_{\perp}=10^{-6}\,m,\,1/p_{\perp}\gtrsim 100\,\text{nm}$. The behaviour at small angles is shown in the insets. The shifts vanish in those regions that are not allowed by the momentum conservation law, they stay roughly the same for other $p^{\prime}_{\perp},p^{\prime}_{z}$ and for ultrarelativistic electrons, $\gamma\gg 1$, though the Cherenkov angle grows.
Figure 5: Dependence of the shift $\Delta t=t_{0}-t_{\text{cl.}}={\bm{l}}_{0}\cdot(\partial_{{\bf p}}+\partial_{{\bf k}})\zeta_{fi}$ from Eq.(16) of the photon arrival time on the refractive index $n$ at the emission angle $\theta=10$ deg. Left: $\beta=0.7,p^{\prime}_{z}=0.9\times\beta$ (a TEM regime; the minimal refractive index for which the Cherenkov condition is met is $n\approx 1.429$, whereas $n=1.45-1.55$ for fused silica in the given spectral range), right: $\beta=0.9999(\gamma=70.7),p^{\prime}_{z}=0.99\times\beta$ (an accelerator regime). The classical Cherenkov angle $\theta_{\text{Ch.cl.}}=\arccos 1/\beta n$ is shown in the insets.
@@
When measuring the quantum shift in the photon arrival time from its classical value, the flash duration $\sigma_{t}(t^{\prime})$ is crucial because the deviations can hardly be discerned when $\sigma_{t}(t^{\prime})\gg|\Delta t|$ occurring for $t^{\prime}\gg t_{d}$ far from the Cherenkov angle. This duration was estimated by Frank as early as in 1956 from the classical considerations [8, 9] to be $\sim 1/\Delta\omega<1\,\text{ps}$ where $\Delta\omega$ is a frequency interval for which the emission takes place. The quantum estimates from the uncertainty principle for the off-resonant AC Stark effect yield the same result [16], $1/\Delta\omega\sim 0.1-100\,\text{fs}$ for $\Delta\omega\sim 10^{-2}-10$ eV. Our quantum model predicts the following flash duration nearby the Cherenkov angle:
@@
because in this case $t_{d}\to\infty$. Clearly, only the electron packet’s size at the target entrance ($t^{\prime}=0$) contributes to the flash duration nearby $\theta_{\text{Ch.cl.}}$. For realistic electrons with $\sigma_{x}(0)\sim 1-100\,\text{nm}$, we find
@@
In Figures 4 and 5 we show that the typical shifts indeed belong to the attosecond range and that the flash duration is larger than the shift, except for small angles. The electron transverse momenta are chosen to be $p_{\perp}=\sigma\sim(10^{-7}-10^{-4})\,m$ because they correspond to the spatial widths $1/p_{\perp}=\sigma_{x}(0)\gtrsim 1\,\text{nm}-1\,\mu\text{m}$, respectively, and we also neglect the spreading. Note that within the pre-wave zone the emission does not take place only at the Cherenkov angle. The sign of the shift swaps between the two kinematic scenarios (Sec.4,5 in Supplementary), which is why the absolute value $|\Delta t|$ is shown. Fixing the detector at certain angles $\theta,\phi$ and the distance $r$, one would see that the photons equally probably arrive either later (time delay) than $t_{\text{cl.}}$ or sooner than that (negative delay), if the electron is not detected. Note that integration of the Wigner function over ${\bf p}^{\prime}$ puts the electron momentum to the definite value ${\bf p}^{\prime}={\bf p}-{\bf k}$ with subsequent integration over ${\bf p}$ with the Gaussian distribution. If one wishes to catch only the shifts with one particular sign, one should detect the photon and the electron in coincidence, which is technically more challenging.
@@
In Figure 5 we highlight that the flash duration $\sigma_{t}$, shown with the green line, can drop below the temporal shift for angles $\theta\lesssim 10-15$ deg and media with small refractive indices $n\sim 1.01-1.5$, however, only with no spreading taken into account as $\sigma_{t}(t^{\prime})$ grows with the time $t^{\prime}$, whereas the shift does not. For the energies of transmission electron microscopes (TEM), $\beta\sim 0.7$, a target made of fused silica with $n=1.458$ in the optical range can do the job, whereas for ultrarelativistic electrons, $\gamma\gg 1$, aerogels with $n\sim 1.01-1.3$ can be employed, which are already used as Cherenkov generators [51, 52, 53]. The use of Brewster-Cherenkov detectors [54] can also come in handy here.
@@
If we now go beyond the model of the electron packet with $\sigma_{\perp}=\sigma_{||}$ and recall that there is Lorentz contraction, $\sigma_{||}=\gamma^{-1}\sigma_{\perp}$, the flash durations can become $\gamma$ times shorter than (18). Although photon spreading can be safely neglected nearby the Cherenkov angle, spreading of the electron before entering the target can significantly increase these numbers. According to the quantum dynamics of the coherence length $\sigma_{x}(t^{\prime})$ – also called the generalized van Cittert-Zernike theorem [41] – when an electron is released from a photo-gun or a field emitter with rms sizes $\sigma_{\perp}\sim\sigma_{||}$ [42] of a few nanometers and an energy up to a few tens of eV, it spreads to some tens of nanometers at the distance of 1 $\mu\text{m}$ and it reaches micrometer sizes already at 1 mm. Clearly, electron acceleration – say, in an RF cavity – mitigates the spreading rate, but a realistic estimate of the packet length $\sigma_{||}$ at the target boundary remains unknown.
@@
In practice, achieving attosecond flash durations requires electrons generating radiation with no spreading before the target, just after the emission from a cathode. A vacuum gap of even a few millimeters after the source would yield micrometer-sized electron packets entering the target, which results in picosecond durations (subpicosecond with acceleration to a few MeV) measurable by streak cameras. State-of-the-art Cherenkov counters at accelerators have picosecond time resolution [53, 55], the subfemtosecond resolution can be achieved at X-ray free-electron lasers [56, 57] or with the frequency-resolved optical gating [58], whereas it is nanoseconds for Cherenkov telescopes in gamma-ray astronomy [21] where spreading effects are notable. Attosecond photon pulses – including twisted photons with orbital angular momentum – are usually obtained through high-harmonic generation in the extreme-ultraviolet and even soft X-ray ranges [59, 60], enabling vortex electrons generation via photoionization [60]. The ChR can be a source of twisted photons [26, 61], also in the soft X-ray range, and shaping the spatial and temporal profiles of the electron wave function offers refined tuning of phase-space profile of the photons.

### v3

@@ Results
and it vanishes for classical emission with no recoil, $\omega/\varepsilon\to 0$. For materials like Al, Si, Be, and Ti, Cherenkov radiation can be observed in the soft X-ray range at the frequencies [47, 48, 46] $\omega\approx 72.5$, $100$, $110$, $453.8$ eV, respectively, which for $\varepsilon\sim(5-20)m$ yields $\Delta\theta_{\infty}<1-2\,\text{deg}$ (cf. Fig.3 c, d).
Between the above points, the time $t_{\text{d}}$ becomes negative – see Fig.3 b – and the Gouy phase $\arctan t^{\prime}/t_{\text{d}}$ changes it sign, as if the electron packet shrinks during the emission. In the classical regime with no recoil, both the points merge and so $t_{\text{d}}$ and $L_{\text{f}}$ turn to infinity at the Cherenkov angle. Indeed, in a vicinity of this angle the spreading time with the recoil kept is
$\displaystyle t_{\text{d}}\Big{|}_{\cos\theta=1/nu_{p}}=\frac{2\varepsilon}{\sigma^{2}}\frac{n^{2}}{1-n^{2}}<0,$ (12)
where $\varepsilon=\gamma m$. We compare this with the spreading time $t_{\text{d}}^{(\text{e, rest})}=m/\sigma^{2}$ of an electron packet, which is at rest on average in vacuum [49]. In the laboratory frame this time is $\gamma$ times larger, which coincides with $|t_{\text{d}}|$ up to the factor $22n^{2}/(n^{2}-1)>$ 2. So, spreading of the photon seems to reverse back in a vicinity of the Cherenkov direction and it is intimately connected with spreading of the electron packet itself.
@@
One can use the peculiar behavior of the diffraction time nearby the Cherenkov angle as a means for detecting the quantum recoil in ChR in UV or soft X-ray range. For that, one needs to measure the size of the photon wave packet at different distances close to the electron path – that is, in the pre-wave zone – and at different polar angles $\theta$ with an angular resolution of at least $0.1$ deg, which is definitely challenging, but not inconceivable.
@@
Dependence of the Wigner function Eq.(5) on the detection time $t$ comes exclusively from the following envelope:
@@
Here, natural duration of the Cherenkov flash is defined by $\sigma_{t}(t^{\prime})$ and $t_{0}={\bm{l}}_{0}\cdot\left({\bf r}+(\partial_{{\bf p}}+\partial_{{\bf k}})\zeta_{\text{fi}}\right)$ is a time instant at which the probability to catch the photon around the point ${\bf r}$ is maximized, ${\bm{l}}_{0}=\left[({\bm{u}}_{p}-{\bm{u}}_{k})\times\left[{\bm{u}}_{k}\times{\bm{u}}_{p}\right]\right]/\left[{\bm{u}}_{p}\times{\bm{u}}_{k}\right]^{2}$. We will call the time $t_{0}$ the mean arrival time.
One can neglect the term with the phase $\zeta_{\text{fi}}$ in ${\bf R}$ Eq.(6) in the wave zone where $|{\bf r}-{\bm{u}}_{p}t|\gg|(\partial_{{\bf p}}+\partial_{{\bf k}})\zeta_{\text{fi}}|$, and then the Wigner function Eq.(5) and the emitted energy seem to be concentrated in a vicinity of the classical electron trajectory [54], ${\bf r}\sim{\bm{u}}_{p}t$. The detector registers a photon in the far field emitted at $t=0,r_{0}=0$ by a classical point-like electron at the time instant
@@
which will be called the classical arrival time. Let us compare this prediction with the above $t_{0}$, derived quantum mechanically. Orienting the $z$ axis along the electron momentum ${\bf p}$, we find ${\bf k}=n\omega{\bm{l}},\ {\bm{l}}=\{\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta\},\ r\equiv{\bf r}\cdot{\bm{l}},\ t_{\text{cl.}}^{(\text{far-f.})}={\bf r}\cdot{\bm{l}}\,n,$ and ${\bm{l}}_{0}=\sin^{-2}\theta\,\left(\left(u_{p}^{-1}-n\cos{\theta}\right){\bf p}/|{\bf p}|+\left(n-u_{p}^{-1}\cos\theta\right){\bm{l}}\right)$. In a vicinity of the Cherenkov angle, we have ${\bm{l}}_{0}\to{\bm{l}}\,n$ and so for $\zeta_{\text{fi}}=0$ we get $t_{0}\to t_{\text{cl.}}^{(\text{far-f.})}={\bf r}\cdot{\bm{l}}\,n$, in accordance with Eq.(15).
@@
However, the phase $\zeta_{\text{fi}}$ cannot be ignored in the formation zone and in the near-field (in the latter case, ${\bf r}\approx{\bm{u}}_{p}t$, so that ${\bf R}\approx(\partial_{{\bf p}}+\partial_{{\bf k}})\zeta_{\text{fi}}$) and this makes the photon arrival time $t_{0}$ different from the classical value $t_{\text{cl.}}^{(\text{far-f.})}$. In this regime, we call $t_{\text{cl.}}={\bf r}\cdot{\bm{l}}_{0}$ for arbitrary emission angles and so the quantum shift is
$\Delta t=t_{0}-t_{\text{cl.}}={\bm{l}}_{0}\cdot(\partial_{{\bf p}}+\partial_{{\bf k}})\zeta_{\text{fi}}.$ (16)
@@
This shift is not necessarily positive and the physical origin of this delay or advance is the electric dipole moment density $\propto e(\partial_{{\bf p}}+\partial_{{\bf k}})\zeta_{\text{fi}}$ induced in medium by the field of the electron. We deal with an analogue of the AC Stark effect [16] with the atoms being off-resonantly polarized by a broadband spectrum $\Delta\omega$ of pseudo-photons. Similarly to the observed time delays – positive and negative – when a laser propagates in a medium [55, 56, 57], here we encounter delays induced by the virtual photons, reemitted as real ones. Classically, one can look at this as if the photon was emitted not from a point-like electron, but from a point shifted laterally to the distance $\Delta\rho\sim\beta\gamma\lambda/2\pi$ from the electron trajectory closer or further from the detector, which is a mean free path of the virtual photon [32, 12]. Numerically $\Delta\rho/c\sim\beta\gamma\lambda/2\pi c=\beta\gamma/\omega\sim 1\,\text{fs}-100\,\text{fs}$ for photons from IR to UV ranges and $\gamma=\varepsilon/m\lesssim 10$.
@@
Figure 4: The quantum shift of the photon arrival time as a function of the emission angle. We take the electron energy typical for a transmission electron microscope: $\varepsilon_{c}\approx 200$ keV, $\beta\approx 0.7$, $p^{\prime}_{\perp}=0.99\times p_{\perp},\,p^{\prime}_{z}=p_{z}-k_{z}\approx 0.98\,m,\,n=1.5$. The electron momentum uncertainty is equal to the transverse momentum: $\sigma=p_{\perp}$. In panel (a): $p_{\perp}=10^{-5}\,m,\,1/p_{\perp}\gtrsim 10\,\text{nm}$; panel (b): $p_{\perp}=10^{-6}\,m,\,1/p_{\perp}\gtrsim 100\,\text{nm}$. The classical Cherenkov flash durations for the emitted photon energies $\omega=10^{-4}m$ (green solid line) and $10^{-7}m$ (brown dashed line) are compared to the quantum shifts (Eq. (16)) for $\omega=10^{-4}m$ (black solid), $10^{-5}m$ (red dashed), $10^{-6}m$ (blue dotted) and $10^{-7}m$ (orange dot-dashed line). The behavior at small angles is shown in the inset figures. The quantum shifts are restricted to the regions allowed by the momentum conservation law and vanish outside of them. They stay roughly the same for other values of $p^{\prime}_{\perp},\,p^{\prime}_{z}$ and for ultrarelativistic electrons, $\gamma\gg 1$, though the Cherenkov angle grows.
Figure 5: The quantum shift of the photon arrival time and the classical Cherenkov angle as a function of the refractive index. Here, the emission angle is chosen to be $\theta=10$ deg and the electron momentum uncertainty is $\sigma=p_{\perp}$. We distinguish a regime of a transmission electron microscope – panel (a): $\beta\approx 0.7\,(\gamma\approx 1)$, and an accelerator regime – panel (b): $\beta=0.9999\,(\gamma\approx 70)$. In both panels (a) and (b) we fix $p^{\prime}_{\perp}=0.99\times p_{\perp},\,p^{\prime}_{z}=p_{z}-k_{z}$, the classical Cherenkov flash duration is shown for $\omega=10^{-5}m$ (green solid line), whereas the quantum shifts are given for $\omega=10^{-5}m$ (black solid line) and $10^{-6}m$ (red dashed line). Panels (c) and (d) demonstrate dependence of the classical Cherenkov angle $\theta_{\text{Ch.cl.}}=\arccos(1/\beta n)$ on the refractive index for two types of dielectric media - fused silica and aerogels. The blue shaded areas in (c) and (d) show the boundaries of the refractive indices due to the frequency dispersion of $n(\omega)$ for fused silica or different types of particular chemical solutions for aerogels. The minimal refractive index for which the Cherenkov condition is met in scenario (a) is $n\approx 1.444$, and $\theta_{\text{Ch.cl.}}=10$ deg corresponds to $n\approx 1.466$ within the range of $n$ from panel (c). Analogously, in scenario (b)$:\theta_{\text{Ch.cl.}}=10$ deg corresponds to $n\approx 1.016$ within the range of $n$ from panel (d).
@@
When measuring the quantum shift in the photon arrival time from its classical value, the flash duration $\sigma_{t}(t^{\prime})$ is crucial because the deviations can hardly be discerned with $\sigma_{t}(t^{\prime})\gg|\Delta t|$ occurring for $t^{\prime}\gg t_{\text{d}}$ far from the Cherenkov angle. This duration was estimated classically by Frank [8, 9] to be $\sim 1/\Delta\omega<1\,\text{ps}$ where $\Delta\omega$ is a frequency interval for which the emission takes place. Quantum estimates from the uncertainty principle yield roughly the same result both in the macroscopic theory of ChR [30] and for the microscopic off-resonant AC Stark effect in single atoms [16], $1/\Delta\omega\sim 0.1-100\,\text{fs}$ for $\Delta\omega\sim 10^{-2}-10$ eV. Our quantum model predicts the following flash duration nearby the Cherenkov angle:
@@
because in this case $t_{\text{d}}\to\infty$. Clearly, only the electron packet’s length at the target entrance ($t^{\prime}=0$) contributes to the flash duration nearby $\theta_{\text{Ch.cl.}}$. For realistic electrons with $\sigma_{x}(0)\sim 1-100\,\text{nm}$, we find
@@
In Figures 4 and 5 we show that the typical shifts indeed belong to the attosecond range and that the flash duration is generally larger than the shift. The electron transverse momenta are chosen to be $p_{\perp}=\sigma\sim(10^{-7}-10^{-4})\,m$ because they correspond to the spatial widths $1/p_{\perp}=\sigma_{x}(0)\gtrsim 1\,\text{nm}-1\,\mu\text{m}$, respectively, and we also neglect the spreading. Note that within the pre-wave zone the emission does not take place only at the Cherenkov angle. The sign of the shift swaps between the two kinematic scenarios (Supplementary Notes 4 and 5), which is why the absolute value $|\Delta t|$ is shown. Fixing the detector at certain angles $\theta,\phi$ and the distance $\bf r$, one would see that the photons equally probably arrive either later (time delay) than $t_{\text{cl.}}$ or sooner than that (an advance or negative delay), if the electron is not detected. Note that integration of the Wigner function over ${\bf p}^{\prime}$ puts the electron momentum to the definite value ${\bf p}^{\prime}={\bf p}-{\bf k}$ with subsequent integration over ${\bf p}$ with the Gaussian distribution. If one wishes to catch only the shifts with one particular sign, one should detect the photon and the electron in coincidence, which is technically more challenging.
@@
From Figures 4 and 5 we conclude that the difference between the classical flash duration $\sigma_{t}$ (green line in both figures) and the temporal shift is minimized for angles $\theta\lesssim 10$ deg and media with small refractive indices $n\sim 1.01-1.5$, whereas the ratio $\sigma_{t}/|\Delta t|\lesssim 2$ for all the angles $\theta<10$ deg. The argument is correct only when spreading is not taken into account because $\sigma_{t}(t^{\prime})$ grows with time $t^{\prime}$ and the shift does not. For larger angles, $\theta\gtrsim 10$ deg, or for $n\gtrsim 2$ the difference between the flash duration (the green line) and the shift increases. When trying to detect the quantum shift for the transmission electron microscope (TEM) energies, $\beta\sim 0.7$, a target made of fused silica with $n\sim 1.44-1.47$ (see, for example, [58, 59]) in the optical range can do the job, whereas for ultrarelativistic electrons, $\gamma\gg 1$, aerogels with $n\sim 1.01-1.30$ can be employed (see, for example, [60, 61]), which are already used as Cherenkov generators [62, 63, 64]. The use of Brewster-Cherenkov detectors [65] can also come in handy here. Note, however, that for the TEMs energies there also be transition radiation generated at a boundary of a real target, which propagates at the angles of $\theta\sim\gamma^{-1}$ and can, therefore, interfere with the Cherenkov signal. Although microscopically this type of radiation has a similar origin as ChR and, therefore, similar temporal shifts are likely to take place, one can start with studying pure Cherenkov emission. One of the alternatives would be to employ the so-called diffraction Cherenkov radiation [46, 12] where an electron moves close to a dielectric target but does not intersect it. In that case, diffraction radiation is generated at the angles $\theta\ll\gamma^{-1}$ and Cherenkov emission can still be well distinguished at $\theta\lesssim 10$ deg.
@@
If we now go beyond the simplified model of the electron packet with $\sigma_{\perp}=\sigma_{||}$ and recall that there is Lorentz contraction, $\sigma_{||}=\gamma^{-1}\sigma_{\perp}$, the flash durations can become $\gamma$ times shorter than Eq.(18). Although photon spreading can be safely neglected nearby the Cherenkov angle, spreading of the electron before entering the target can significantly increase these numbers. According to the quantum dynamics of the electron coherence length $\sigma_{x}(t^{\prime})$ – also called the generalized van Cittert-Zernike theorem [49] – when an electron is released from a photo-gun or a field emitter with rms sizes $\sigma_{\perp}\sim\sigma_{||}$ [50] of a few nanometers and an energy up to a few tens of eV, it spreads to some tens of nanometers at the distance of 1 $\mu\text{m}$ and it reaches micrometer sizes already at 1 mm from the source. Clearly, electron acceleration – say, in an RF cavity – mitigates the spreading rate, but a realistic estimate of the packet length $\sigma_{||}$ at the target boundary remains unknown and it is not usually controlled. Therefore, measuring the flash duration in a vicinity of the ChR angle, one can retrieve the length of the electron packet at the target entrance, a complementary approach to interferometry that is reminiscent of bunch length measurements via classical coherence [14].
@@
In practice, achieving attosecond flash durations requires nanometer-sized electron packets generating radiation with no spreading before the target, just after the emission from a cathode. A vacuum gap of even a few millimeters after the source and before the target would yield micrometer-sized electron packets entering the target, which results in picosecond flash durations or in subpicosecond ones with acceleration to a few MeV in the gap due to the above factor $\gamma^{-1}$, measurable by streak cameras. State-of-the-art Cherenkov counters at accelerators have picosecond time resolution [64, 66], the subfemtosecond resolution can be achieved at X-ray free-electron lasers [67, 68] or with the frequency-resolved optical gating [69], whereas it is nanoseconds for Cherenkov telescopes in gamma-ray astronomy [21] where spreading effects are notable. Attosecond photon pulses – including twisted photons with orbital angular momentum – are usually obtained through high-harmonic generation in the extreme-ultraviolet and even soft X-ray ranges [70, 71], enabling vortex electrons generation via photoionization [71]. The ChR can be a source of twisted photons [26, 31], also in the soft X-ray range, and shaping the spatial and temporal profiles of the electron wave function offers refined tuning of phase-space profile of the photons. In particular, it seems feasible to generate photon wave packets with the given spatial profile, vorticity, and the energies up to the soft X-ray range with the attosecond and even sub-attosecond durations.
Conclusions